OpenStudy (anonymous):

The school cafeteria sells two kinds of wraps: vegetarian and chicken. The vegetarian wrap costs $1.00 and the chicken wrap costs $1.80. Today they made $98.80 from the 70 wraps sold. How many of the wraps sold were vegetarian? A. 36 wraps B. 37 wraps C. 30 wraps D. 34 wraps A, B, C, or D? Please explain (:

OpenStudy (anonymous):

34 should be correct . pick answer D

OpenStudy (anonymous):

Thanks much :D

OpenStudy (anonymous):

you are welcome :) good luck

OpenStudy (ajprincess):

Let the number of vegetarian wraps be v and the number of chicken wraps be c. v+c=70-(1) 1.0v+1.8c=98.8 10v+18c=988-(2) (1)*10=>10v+10c=700-(3) (2)-(3)=>8c=288 c=288/8 c=36 v+36=70 v=70-36 =34